# 2-Cats: 2D Copula Approximating Transforms

### v3

@@ 3.2 Proofs
Now, considering a bivariate CDF, for any given point in its domain (e.g., $t_{v}(u),t_{u}(v)$), the CDF represents the accumulated probability up to that point. Let the origin be $(t_{v}(0),t_{u}(0))$. As either $t_{v}(u)$, $t_{u}(v)$, or both values increase, the CDF values can only increase or remain constant. Therefore, the volume under the CDF surface will always be positive. This proves P3. ∎
2-Cats is a Copula with Positive Density and Derivatives that Capture CDFs.
The second derivative of 2-Cats is a probability density function, the first derivative is a CDF.
Proof.
As a consequence of P[1-3]. These three properties define a Copula [Nelsen, 2006]. Thus, the first derivative of 2-Cats is the corresponding conditional CDF, and the second derivative of the model is associated density function [Nelsen, 2006, Theorem 2.2.7 and Theorem 2.2.8]. ∎
A mixture of 2-Cats is a Universal Approximator for Copulas.
Mixture Distributions Nguyen et al. [2020] can approximate any other probability distributions. If we define a model that is a weighted mixture of 2-Cats models, it can approximate any other Copula. Details on how to build and train such a model are on Appendix B
4 Training 2-Cats Models
The inputs of our model are comprised of the set $\mathcal{D}=\{u_{i},v_{i}\}$, where $i\in[1,n=|\mathcal{D}|]$. We make use of subscripts to index data points, dimensions, as well as random variables (RVs). For instance, $x_{i,j}$ is the $j$-th dimension of the $i$-th data point. When working with RVs, we do not need data point indexes, thus we omit them and leave subscripts for dimensions only. $X_{j}$ is the RV for the $j$-th dimension.
4.1 Overall Loss
@@
$L_{\bm{\theta}}(\mathcal{D})=w^{C}L^{C}_{\bm{\theta}}(\mathcal{D})+w^{dC}L^{dC}_{\bm{\theta}}(\mathcal{D})+w^{c}L^{c}_{\bm{\theta}}(\mathcal{D})$
@@
The first component, $L^{C}_{\bm{\theta}}(\mathcal{D})$, stimulates the copula to closely resemble the empirical cumulative distribution function derived from the observed data. In this way, our model learns to capture the essential characteristics of the data. The second component, $L^{dC}_{\bm{\theta}}(\mathcal{D})$, imposes penalties on any disparities between the fitted copula’s first-order derivatives and the data-based estimates. The third component, $L^{c}_{\bm{\theta}}(\mathcal{D})$, focuses on the copula’s second-order derivative, which is linked to the probability density function of the data, and evaluates its proximity to the empirical likelihood. Incorporating this aspect in our loss function enhances the model’s ability to capture the data’s distribution.
While obtaining the first component is relatively straightforward, the other two components necessitate some original work. We explain each of them in turn.
4.2 Base ECDF Loss
Let $F(\mathbf{x})=\Pr\big{[}X_{1}\leq x_{1},X_{2}<x_{2},\dots,X_{d}\leq x_{d}\big{]}$ be the $d$-variate CDF for some random vector $\mathbf{x}$ and $\widehat{F^{n}}(\mathbf{x})$ be the empirical cumulative distribution function (ECDF). This ECDF will be estimated from a sample of size $n$.
@@ 4.2 Base ECDF Loss
We can explore these results to define the relationship between our model, the ECDF, and the true CDF. Given that the CDF $F(x_{1},x_{2})$ and the Copula $C(u,v)$ evaluate to the same value (i.e., $u$ and $v$ are the inverse of the marginal CDFs of $x_{1}$ and $x_{2}$ respectively):
$C(u,$ $v)=F(F_{x_{1}}^{-1}(u),F_{x_{2}}^{-1}(v))=F(x_{1},x_{2})\approx$
$\widehat{F^{n}}(x_{1},x_{2})\approx H_{\theta}(u,v)=H_{\theta}(\widehat{F_{x_{1}}}(x_{1}),\widehat{F_{x_{2}}}(x_{2})).$
@@
Considering these definitions, we can define the loss functions based on the ECDF and the model’s output:
$L^{C}_{\bm{\theta}}(\mathcal{D})=\frac{1}{n}\sum_{u_{i},v_{i}\in\mathcal{D}}\big{(}H_{\bm{\theta}}(u_{i},v_{i})-\widehat{F^{n}}(x_{i,1},x_{i,2})\big{)}^{2}.$
Figure 1: Models vs Derivatives. Here, we compare sample data from the simple $y=sin(x)$ process. Data is sampled in three settings, a “full” dataset with a larger range (left), a noisy dataset (middle), and a limited view (right). NN models (lines) approximate the data points in every case. However, the derivatives/integrals of the model do not approximate function derivatives (last row). Integrals (first row) are approximated up to a constant Liu [2023].
4.3 Copula First Derivative Loss
The MSE above helps approximate the bivariate CDF of the Copula. However, we have no guarantees that it will approximate the derivatives of the Copula. To understand this argument, we take some time to revise the universal approximation properties of NNs [Cybenko, 1989, Hornik et al., 1989, Lu and Lu, 2020]. We also shift our focus from Copulas for this argument. It is expected to state that the universal approximation theorem (UAT) guarantees that functions are approximated with NNs. However, given that when learning models from training data, it would be more correct to say that function evaluations at training points are approximated.
Consider a simple 4-Layer Relu NN with layers of 128 neurons. This will be our model. To avoid overfitting, this model will be optimized with a validation set. The problem with relying only on the UAT is shown in Figure 1. The figure shows a ground truth data-generating process of $y=\sin(x)+\eta$, where $\eta$ is some noise term. On the first column, when $\eta=0$, the model will approximate the data generating function (middle row), its derivative (bottom row), and integral (top row). When $\eta\sim N(0,0.5)$, the model will approximate the data points, but not the true derivatives. Integrals are approximated up to an asymptotic constant (see Appendix C and  [Liu, 2023]); however, we still need to control for this constant (note the growing deviation on the integral on the top plot of the middle row).
@@ 4.3 Copula First Derivative Loss
Returning to Copulas, a natural step for the first derivatives would be to define a mean squared error (or similar) loss towards these derivatives. The issue is that we need to define empirical estimates from data for both derivatives: $\frac{\partial C(u,v)}{\partial u}$ and $\frac{dC(u,v)}{\partial v}$. To estimate such derivatives, we could have employed conditional density estimators Papamakarios et al. [2017], Sohn et al. [2015]. However, we explore the underlying properties of a Copula to present an empirical estimate for such derivatives.
For a 2d Copula, it is well known that the first derivative of $C$ has the form [Nelsen, 2006]:
$\frac{\partial C(u,v)}{\partial u}=\Pr\big{[}V\leq v\mid U=u\big{]}.$
@@
The issue with estimating this function from data is that we cannot filter our dataset to condition on $U=u$ (we are working with data where $u\in\mathbb{I}$, a real number). However, using Bayes rule we can rewrite this equation.
First, let us rewrite the above equation in terms of density functions. That is, if $c_{u}(v)=F(V\leq v\mid U=u\big{)}=\Pr\big{[}V\leq v\mid U=u\big{]}$ is a cumulative function, $c(v)$ and $c(u)$ are the marginal copula densities (uniform by definition). Using Bayes rule, we have that:
$\frac{\partial C(u,v)}{\partial u}$ $=c_{u}(v)=\frac{c\big{(}U=u\mid V\leq v\big{)}F\big{(}V\leq v\big{)}}{c\big{(}u\big{)}}$
$=v\,c\big{(}U=u\mid V\leq v\big{)}.$
@@
Where the last identity above comes from the fact that for Copulas, the marginal distributions $U$ and $V$ are uniform, leading to $c(u)=c(v)=1$. Also, we shall have that $F\big{(}V\leq v\big{)}=\Pr\big{[}V\leq v\big{]}=v$ and $F\big{(}U\leq u\big{)}\Pr\big{[}U\leq u\big{]}=u$ [Nelsen, 2006]. What is left now is to estimate the density of $u$ for values of $V\leq v$, i.e.: $c\big{(}U=u\mid V\leq v\big{)}$.
@@
To do so, let us define $\frac{\widehat{\partial C^{n}}(u,v)}{\partial u}\approx c\big{(}U=u\mid V\leq v\big{)}$ as an empirical estimate of such a density using $n$ data points. We employ the widely used Kernel Density Estimation (KDE) to estimate this function. To do so, we can arrange our dataset as a table of two columns, where each row contains the pairs of $u,v$ (the columns). For efficiency, we create views of this table that are sorted on column $u$ or column $v$.
When we iterate this table sorted on $v$, finding the points where $V\leq v$ is trivial, as these are simply the previously iterated rows. If we perform KDE on the column $u$ for these points, we shall have an empirical estimate of the density: $c\big{(}U=u\mid V\leq v\big{)}$. By simply multiplying this estimate with $v$, we have our empirical estimation of the partial derivative with regards to $u$, that is: $\frac{\widehat{\partial C^{n}}(u,v)}{\partial u}$. We can perform a similar approach to estimate the partial derivative on $v$: $\frac{\widehat{\partial C^{n}}(u,v)}{\partial v}$. Now, our loss term is:
$(\displaystyle L^{dC}_{\bm{\theta}}(\mathcal{D})=\frac{1}{2n}\sum_{u_{i},v_{i}\in\mathcal{D}}\big{(}$$(\frac{\partial H_{\bm{\theta}}(u_{i},v_{i})}{\partial u}-\frac{\widehat{\partial C^{n}}(u_{i},v_{i})}{\partial u})^{2}+$
$(\frac{\partial H_{\bm{\theta}}(u_{i},v_{i})}{\partial v}-\frac{\widehat{\partial C^{n}}(u_{i},v_{i})}{\partial v})^{2}\big{)}.$
4.4 Pseudo-Likelihood Loss
[…]
$L^{c}_{\bm{\theta}}(\mathcal{D})=-\frac{1}{n}\sum_{u_{i},v_{i}\in\mathcal{D}}\log(\frac{\partial^{2}H_{\bm{\theta}}(u_{i},v_{i})}{\partial u\partial v}).$
The literature refers to this loss as pseudo-log-likelihood [Genest et al., 1995, Genest and Favre, 2007]. Its asymptotic properties have been well characterized [Genest and Favre, 2007], and it is commonly employed in Copula models. It is referred to as a pseudo-likelihood because it is a likelihood of the transformed dataset (in Copula form).
@@ 4.4 Pseudo-Likelihood Loss
We point out to the reader that it is essential to understand why the three losses are needed. It might seem that simply minimizing this loss (or maximizing this log-likelihood) would be sufficient. This is not so for two reasons. Firstly, as depicted in Figure 1, the integrals of a model will not necessarily follow the integrals of the underlying function. We can show that integrals of NN are also approximations up to an asymptotical constant Liu [2023], but we still need to ensure that this constant is acceptable (see also Appendix C).
Secondly, naively maximizing this single, $L^{c}_{\bm{\theta}}(\mathcal{D})$, the loss would lead to NN that always outputs large (possibly $\infty$) values on a training set.
Our other losses, $L^{C}_{\bm{\theta}}(\mathcal{D})$ and $L^{dC}_{\bm{\theta}}(\mathcal{D})$ will thus act as regularizers for this third loss to make it so that $h_{\bm{\theta}}(u,v)$ acts like a valid likelihood.

### v4

@@ 3 On the Desiderata of a Copula and 2-Cats Properties
Considering a bivariate CDF, for any given point in its domain (e.g., $t_{v}(u),t_{u}(v)$), the CDF represents the accumulated probability up to that point. Let the origin be $(t_{v}(0),t_{u}(0))$. As either $t_{v}(u)$, $t_{u}(v)$, or both values increase, the CDF values can only increase or remain constant. Therefore, the volume under the CDF surface will always be positive. Now, consider the following Corollary of Eq 2. ∎
Corollary: The second derivative of 2-Cats is a pseudo-likelihood..
Let $F(x_{1},x_{2})=\Pr[X_{1}\leq x_{1},X_{2}\leq x_{2}]$ be the bivariate CDF associated with RVs $X_{1}$ and $X_{2}$. With $u=F_{X_{1}}(x_{1})$ and $v=F_{X_{2}}(x_{2})$ being the marginal CDFs, we reach that: $F(x_{1},x_{2})=\Pr\big{[}U\leq u,V\leq v\big{]}$. By Eq (2), $\Pr\big{[}U\leq u,V\leq v\big{]}=H_{\bm{\theta}}(u,v)=H_{\bm{\theta}}(F_{X_{1}}(x_{1}),F_{X_{2}}(x_{2}))$. 2-Cats is also capturing the same bivariate CDF of $C$. The bivariate density, $f(x_{1},x_{2})=\frac{\partial^{2}F(x_{1},x_{2})}{\partial x_{1}\,\partial x_{2}}$, thus is:
$f(x_{1},x_{2})=\frac{\partial^{2}H_{\bm{\theta}}(F_{X_{1}}(x_{1}),F_{X_{2}}(x_{2}))}{{\partial x_{1}\,\partial x_{2}}}=f_{X_{1}}(x_{1})\,f_{X_{2}}(x_{2})\frac{\partial^{2}H_{\bm{\theta}}(u,v)}{{\partial u\,\partial v}}$ (3)
The above equation is solved via the Chain rule. A consequence of it, is that the likelihood $f(x_{1},x_{2})$, is proportional to $h_{\bm{\theta}}(u,v)=\frac{\partial^{2}H_{\bm{\theta}}(u,v)}{{\partial u\,\partial v}}$. This term is known as a pseudo-likelihood [23, 22].∎
Proof P2..
Given that $f(x_{1},x_{2})\geq 0$ by definition, and also $f_{X_{1}}(x_{1})\geq 0$ and $f_{X_{2}}(x_{2})\geq 0$: $h_{\bm{\theta}}(u,v)\geq 0$. The volume being the double integral of $h_{\bm{\theta}}(u,v)$ is always positive or zero. ∎
Property P3.
$H_{\bm{\theta}}(u,0)=0$, $H_{\bm{\theta}}(0,v)=0$, $H_{\bm{\theta}}(u,1)\approx u$, and $H_{\bm{\theta}}(1,v)\approx v$. Notice that this property is a relaxation of the one presented in Section 1.
This last property of a Copula is the one that guarantees that the Copula will have uniform marginals. In particular, the terms: $C(u,1)=u$ and $C(1,v)=v$ are of utmost importance for sampling. To grasp this, consider some valid Copula $C$. Here, $C(u,1)=\Pr[U\leq u,V\leq 1]=\Pr[U\leq u]=u$, which is the CDF of the Uniform distribution. A similar argument exists for $v$.
2-Cats does not provably meet P3 as it does P1 and P2. Nevertheless, we prove that: $H_{\bm{\theta}}(u,0)=0$ and $H_{\bm{\theta}}(0,v)=0$, while approximating $H_{\bm{\theta}}(u,1)\approx u$, and $H_{\bm{\theta}}(1,v)\approx v$.
Lemma P3.1.
$H_{\bm{\theta}}(u,0)=0$ and $H_{\bm{\theta}}(0,v)=0$
Proof..
Consider the value of $H_{\bm{\theta}}(u,0)$ when $v=0$. As $t_{u}(0)=0$, we have $z(t_{u}(0))=z(0)=\lim_{x\rightarrow 0}z(x)=-\infty$ and hence $H_{\bm{\theta}}(u,0)=G(z_{u},-\infty)$. Given that $G$ is a bivariate CDF, $G(z_{u},-\infty)=\lim_{w\rightarrow-\infty}G(z_{u},w)=0$ for any $z_{u}$. The same proof exists for $H_{\bm{\theta}}(0,v)$. ∎
Conjecture P3.2.
$H_{\bm{\theta}}(u,1)\approx u$ and $H_{\bm{\theta}}(1,v)\approx v$.
Meeting P3.2 via Lagrangian..
To meet this relaxed property, we propose the following constrained optimization. Let the inputs of our model be comprised of the set $\mathcal{D}=\{u_{i},v_{i}\}$, where $i\in[1,n=|\mathcal{D}|]$. Moreover, let $L_{\bm{\theta}}(D)$ be some loss term used to optimize our model (see the next section). Our optimization for 2-Cats will focus on the following constrained problem:
$\arg\,\min_{\bm{\theta}}L_{\bm{\theta}}(D)\quad\textrm{s.t.}\quad\forall u\in[0,1]:H_{\bm{\theta}}(u,1)-u=0\,\text{and}\,\forall v\in[0,1]:H_{\bm{\theta}}(1,v)-v=0.$ (4)
Although it is trivial to understand why this optimization meets P3.2, performing such an optimization presents several challenges. The first challenge is how to model constraints. A natural first choice is to consider the square of constraints. However, squared constraint optimization is an ill-posed optimization problem [8, Chapter 2.1], and  [54, Section 2.1]. Secondly, our model may correctly estimate that $H_{\bm{\theta}}(u,1)=u$ (or similarly that $H_{\bm{\theta}}(1,v)=v$), but this will not necessarily ensure that the partial derivatives of $H_{\bm{\theta}}(u,1)$ and $H_{\bm{\theta}}(1,v)$ are accurate (see Appendix A.3). Such derivatives must equal one for these inputs as the Copula marginals are distributed according to an $Uniform(0,1)$. Thirdly, evaluating these constraints over the entire domain $u\in[0,1]$ and $v\in[0,1]$ is impossible. The final challenge is how to optimize constraints. Here, a natural approach is to consider Lagrangian multipliers [54, 8, 69]. However, in such cases, the solution to the optimization problem will lie on a saddle point of the loss surface, and gradient methods (used in NNs and 2-Cats) do not converge on saddle points (Section 3.1 of [54] provides a simple intuition on why this is so).
We tackle the first and second challenges by considering the derivative of the square of the above constraints, i.e: $\frac{\partial(H_{\bm{\theta}}(u,1)-u)^{2}}{\partial u}$ and $\frac{\partial(H_{\bm{\theta}}(1,v)-v)^{2}}{\partial v}$, which via the chain rule will be defined as follows:
$r^{1}_{\bm{\theta}}(u)=2\,\big{(}H_{\bm{\theta}}(u,1)-u\big{)}\big{(}\frac{\partial H_{\bm{\theta}}(u,1)}{\partial u}-1\big{)}\,\text{and}\,r^{2}_{\bm{\theta}}(v)=2\,\big{(}H_{\bm{\theta}}(1,v)-v\big{)}\big{(}\frac{\partial H_{\bm{\theta}}(1,v)}{\partial v}-1\big{)}.$ (5)
[…]
We evaluated the constraints on our training set to tackle the third challenge. This leads to the optimization stated below. $r_{\bm{\theta}}(D)$ are the constraints on the training data. From the arguments above, we have that: $r_{\bm{\theta}}(D)\geq 0$ and $r_{\bm{\theta}}(D)=0$ if and only if the constraints are met.
$\hat{\bm{\theta}}=\arg\,\min_{\bm{\theta}}L_{\bm{\theta}}(D)+r_{\bm{\theta}}(D),\,\,\text{where: }\,\,r_{\bm{\theta}}(D)=\sum_{i=1}^{n}r^{1}_{\bm{\theta}}(u_{i})+\sum_{i=1}^{n}r^{2}_{\bm{\theta}}(v_{i}).$ (6)
As in [7], to solve the above problem, we treat $r_{\bm{\theta}}(D)$ as a barrier and optimize as follows:
  1. 1.
    Let $\lambda\in(0,1]$ and $\alpha\in(0,1)$ be hyper-parameters (e.g., we set $\lambda=1$ and $\alpha=0.95$). Also, let $\bm{\theta}^{(0)}$ be the initial model parameters (initialized as in [34] – aka Lecun Normal).
  2. 2.
    For every training iteration $t\in[1,T]$:
    1. (a)
      $\bm{\theta}^{(t)}=\text{stochatic gradient step}\big{(}\lambda\,L_{\bm{\theta}^{(t-1)}}(D)+r_{\bm{\theta}^{(t-1)}}(D)\big{)}$ via ADAM [33]
    2. (b)
      $\lambda=\alpha\,\lambda$
  3. 3.
    Set our model as $\hat{\bm{\theta}}=\bm{\theta}^{(T)}$
As $t\rightarrow\infty$, we have that $\lambda\rightarrow 0$, leading our optimization to focus on the constraints. Even though we may not start our optimization in a feasible solution to $r_{\bm{\theta}^{(t-1)}}(D)$ [7], the fact that our loss surface is fully differentiable allows the optimizer to reach such a solution if not stuck on some local minima/saddle. Nevertheless, we align with the empirical evidence that in overparametrized models, such as deep NNs, local minima are rare (see Chapter 6.7 – The Optimality of Backpropagation – of [6]) and that such stochastic optimizers, such as ADAM, are suitable for escaping saddle points [6]. Moreover, our initialization, $\bm{\theta}^{(0)}$, is suitable for satisfying regularizers [34].
For comparison, we shall present results with and without these Lagrangian terms. ∎
4 Sobolev Losses for 2-Cats Models
@@
$L_{\bm{\theta}}(\mathcal{D})=w^{C}L^{C}_{\bm{\theta}}(\mathcal{D})+w^{dC}L^{dC}_{\bm{\theta}}(\mathcal{D})+w^{c}L^{c}_{\bm{\theta}}(\mathcal{D})$ (7)
@@
The first component, $L^{C}_{\bm{\theta}}(\mathcal{D})$, stimulates the copula to closely resemble the empirical cumulative distribution function derived from the observed data. In this way, our model learns to capture the essential characteristics of the data. The second component, $L^{dC}_{\bm{\theta}}(\mathcal{D})$, imposes penalties on any disparities between the fitted copula’s first-order derivatives and the data-based estimates. A Copula’s first derivative is essential for sampling (see Appendix B) and Vine methods. The third component, $L^{c}_{\bm{\theta}}(\mathcal{D})$, focuses on the copula’s second-order derivative, which is linked to the probability density function of the data, and evaluates its proximity to the empirical likelihood. Incorporating this aspect in our loss function enhances the model’s ability to capture the data’s distribution.
From the arguments above, all three terms play a role in Copula modeling. While obtaining the first component is relatively straightforward, the other two components necessitate some original work. We explain each of them in turn. Before continuing, we point out that Appendix A provides an example of how NNs behave when estimating derivatives and integrals of functions.
@@
We can explore these results to define the relationship between our model, the ECDF, and the true CDF. Being $\hat{\bm{\theta}}$ the estimated parameters and given that the CDF $F(x_{1},x_{2})$ and the Copula $C(u,v)$ evaluate to the same value (i.e., $u$ and $v$ are the inverse of the marginal CDFs of $x_{1}$ and $x_{2}$ respectively):
$C(u,$ $v)=F(F_{X_{1}}^{-1}(u),F_{X_{2}}^{-1}(v))=F(x_{1},x_{2})\approx\widehat{F^{n}}(x_{1},x_{2})\approx H_{\hat{\bm{\theta}}}(u,v)=H_{\hat{\bm{\theta}}}(\widehat{F_{X_{1}}}(x_{1}),\widehat{F_{X_{2}}}(x_{2})).$
@@
Considering these definitions, we can define the loss functions based on the ECDF and the output:
$L^{C}_{\bm{\theta}}(\mathcal{D})=\frac{1}{n}\sum_{u_{i},v_{i}\in\mathcal{D}}\big{(}H_{\bm{\theta}}(u_{i},v_{i})-\widehat{F^{n}}(x_{i,1},x_{i,2})\big{)}^{2}.$ (8)
@@
Now for our second term, $L^{dC}_{\bm{\theta}}(\mathcal{D})$, a natural step for the first derivatives would be to define a mean squared error (or similar) loss towards these derivatives. The issue is that we need to define empirical estimates from data for both derivatives: $\frac{\partial C(u,v)}{\partial u}$ and $\frac{dC(u,v)}{\partial v}$. To estimate such derivatives, we could have employed conditional density estimators [53, 63, 26]. Nevertheless, methods as [53, 63] are also deep NNs that have the drawback of requiring extra parameters to be estimated. Even classical methods such as [26] suffer from this issue. As a minor contribution, we explore the underlying properties of a Copula to present an empirical approximation for such derivatives.
For a 2d Copula, the first derivative of $C$ has the form [50]: $\frac{\partial C(u,v)}{\partial u}=\Pr\big{[}V\leq v\mid U=u\big{]}.$
@@
The issue with estimating this function from data is that we cannot filter our dataset to condition on $U=u$ (we are working with data where $u\in\mathbb{I}$, a real number). However, using Bayes rule we can rewrite this equation. First, let us rewrite the above equation in terms of density functions. That is, if $c_{u}(v)=F(V\leq v\mid U=u\big{)}=\Pr\big{[}V\leq v\mid U=u\big{]}$ is a cumulative function, $c(v)$ and $c(u)$ are the marginal copula densities (uniform by definition). Using Bayes rule, we have that:
$\frac{\partial C(u,v)}{\partial u}$ $=c_{u}(v)=\frac{c\big{(}U=u\mid V\leq v\big{)}F\big{(}V\leq v\big{)}}{c\big{(}u\big{)}}=v\,c\big{(}U=u\mid V\leq v\big{)}.$ (9)
@@
Where the last identity above comes from the fact that for Copulas, the marginal distributions $U$ and $V$ are uniform, leading to $c(u)=c(v)=1$. Also, we shall have that $F\big{(}V\leq v\big{)}=\Pr\big{[}V\leq v\big{]}=v$ and $F\big{(}U\leq u\big{)}\Pr\big{[}U\leq u\big{]}=u$ [50]. Now, estimate $c\big{(}U=u\mid V\leq v\big{)}$.
@@
To do so, let us define $\frac{\widehat{\partial C^{n}}(u,v)}{\partial u}\approx c\big{(}U=u\mid V\leq v\big{)}$ as an empirical estimate of such a density using $n$ data points. We employ the widely used Kernel Density Estimation (KDE) to estimate this function. A fast algorithm for our estimation will work as follows: we arrange our dataset as a table of two columns, where each row contains the pairs of $u,v$ (the columns). For efficiency, we create views of this table sorted on column $u$ or column $v$. When we iterate this table sorted on $v$, finding the points where $V\leq v$ is trivial, as these are simply the previously iterated rows. If we perform KDE on the column $u$ for these points, we shall have an empirical estimate of the density: $c\big{(}U=u\mid V\leq v\big{)}$. By simply multiplying this estimate with $v$, we have our empirical estimation of the partial derivative with regards to $u$, that is: $\frac{\widehat{\partial C^{n}}(u,v)}{\partial u}$. Similarly, we can estimate $\frac{\widehat{\partial C^{n}}(u,v)}{\partial v}$. Now, our loss term is:
$(\displaystyle L^{dC}_{\bm{\theta}}(\mathcal{D})=\frac{1}{2n}\sum_{u_{i},v_{i}\in\mathcal{D}}\big{(}$$(\frac{\partial H_{\bm{\theta}}(u_{i},v_{i})}{\partial u}-\frac{\widehat{\partial C^{n}}(u_{i},v_{i})}{\partial u})^{2}+(\frac{\partial H_{\bm{\theta}}(u_{i},v_{i})}{\partial v}-\frac{\widehat{\partial C^{n}}(u_{i},v_{i})}{\partial v})^{2}\big{)}.$ (10)
Finally, we focus on the last part of our loss, the a pseudo-likelihood:
$L^{c}_{\bm{\theta}}(\mathcal{D})=-\frac{1}{n}\sum_{u_{i},v_{i}\in\mathcal{D}}\log(\frac{\partial^{2}H_{\bm{\theta}}(u_{i},v_{i})}{\partial u\partial v}).$ (11)
@@
It is essential to understand why the three losses are needed. It might seem that simply minimizing this loss would be sufficient. However, we can show that integrals of NN are also approximations up to an asymptotical constant [39], but we still need to ensure that this constant is acceptable (see Appendix A.2). Moreover, Appendix C presents an ablation study on the impact of all three losses.